# Artificial Intelligence in Creative Industries: Advances Prior to 2025

### v2

@@ 3.7.2 Performance and main challenges
It should be noted that due to the lack of standard test conditions and limited model generalization within many existing image and video quality assessment models, these methods are typically trained and benchmarked with different databases and intra-database cross-validation is often performed. This results in different evaluation results and conclusions. To enable a fair and meaningful comparison, various challenges and contests have been held for visual quality assessment. The Sixth Challenge on Learned Image Compression (CLIC) [295] associated with the Data Compression Conference 2024 is one of the latest examples which includes two quality assessment tracks for image and video compression. The best performer in the video quality assessment track achieves a Spearman Ranking Correlation Coefficient value of 0.825 [227]. Other notable challenges include the IEEE/CVF WACV 2023 HDR VQA Grand Challenge and the Video Super-Resolution Quality Assessment Challenge in ECCV 2024, which focus on high dynamic rage and super-resolved content, respectively.
Although significant progress has been made in the past few years in visual quality assessment, including new models and training methodology, challenges still exist, including limited model generalization and high computational complexity. Another important application of quality metrics is to be used as a loss function for different image and video processing methods, which is expected to produce results with better visual quality. This requires further capability and robustness for these quality models. All of these issues need to be addressed in future work.
4 Future Challenges for AI in the Creative Sector
LLMs enable artists to articulate their ideas through text prompts, facilitating a top-down creative process. This approach starts with a high-level conceptualization and planning, allowing artists to frame an overarching idea, theme, or goal that guides the creative process before diving into specific details or implementation tasks. This method contrasts with the bottom-up approach, where creativity begins with detailed elements and builds up to a broader structure or idea. This means the effectiveness of current AI technologies may vary among artists, and achieving desirable outcomes from AI heavily depends on the users’ ability to provide accurate text inputs. Moreover, recent technologies generate digital content from a ‘random’ seed, meaning that with the same text prompts, the output might not be the same. This makes editing those outputs to better fit the precise thoughts or imagination of artists still challenging.
Originally not directly designed for creative industries, coding and debugging tools [375] help software developers expedite their work, including tasks within creative sectors. Speaking at the World Government Summit in Dubai in 2024484848https://blogs.nvidia.com/blog/world-governments-summit/, Nvidia CEO Jensen Huang argued that with rapid advancements in AI, learning to code might become less critical for those entering the tech sector. He foresees a future where advancements in computing might make traditional programming obsolete, thus automating tasks and boosting productivity, especially for artists who lack programming expertise. While this view isn’t universally accepted, it clearly offers potential benefits for artists navigating an evolving tech landscape. However, AI coding for creative artists may need specific requirements to serve unstructural creativity.
Furthermore, leading AI organizations like OpenAI, DeepMind, and Anthropic are proponents of the potential for artificial general intelligence (AGI), an advanced form of AI that could achieve human or even superhuman levels of intelligence. Such AI systems could eventually exceed human intelligence, potentially benefiting humanity as a whole. However, solving such complex problems requires not only advanced computation and innovative algorithms but also a deep understanding of how the human brain functions.
@@ 4.1 Ethical Issues, Fakes and Bias
Advancements in AI, exemplified by models like Sora and Gemini 1.5 Pro, provoke ethical concerns and societal implications. These models, capable of generating realistic content, escalate the risk of misuse, including deepfakes and misinformation. Rapid AI development raises questions about job displacement and the delicate balance between automation and human involvement. Ensuring AI augments rather than undermines human efforts poses a significant challenge for developers and policymakers. In a related context, the artist Miles Astray demonstrated that even authentic photographs could be mistaken for AI-generated images. His real photograph titled ‘F L A M I N G O N E’ won both the jury’s award and the people’s choice award in the AI category of the 1839 Awards. He aimed to highlight the ethical dilemmas inherent in AI, suggesting that the benefits of discussing AI’s ethical implications could surpass the ethical concerns related to viewer deception [^49] https://www.milesastray.com/news/newsflash-reclaiming-the-brain.
The uncanny valley effect further complicates matters, challenging perceptions of authenticity. Democratizing AI tools presents opportunities but necessitates robust regulatory frameworks to safeguard privacy. Deepfake technology, while having potential positive applications like in filmmaking, also raises significant concerns about spreading misinformation and other malicious uses. Efforts to detect and combat deepfakes include developing algorithms and tools to identify manipulated media, promoting media literacy, and implementing policies to address the ethical and legal implications.
@@
Unified Concept Editing [55] has been proposed for image generation in digital mediums. This method aims to ensure diverse representation, reducing gender and racial biases, and producing safe content. Additionally, hallucination in generative AI refers to models generating outputs that are not faithful representations of reality but instead contain imagined or unrealistic elements, which may be due to limitations in the training data, biases in the model architecture, or imperfections in the optimization process. The hallucinations is one of the issues that the UK government’s viewpoint on LLMs highlights ongoing issues [1]. Other issues include biases, regurgitation of private data, difficulties with multi-step tasks, and challenges in interpreting black-box processes.
@@
Governments have increasingly expressed concerns about the challenges and key uncertainties that generative AI technologies pose to rightsholders and human creativity [2]. Generative AI presents substantial legal challenges, including the copyright status of AI-generated works and the ownership of intellectual property and copyrights for the datasets used in training AI models, which typically originate from various sources. There are different views from artists on using AI and copyrights. For example, the track “Heart on My Sleeve,” penned by an as yet unidentified human author, featured AI-generated vocals that replicated the voices of Drake and The Weeknd. Released independently on April 4, 2023, it was accessible on various streaming platforms including Apple Music, Spotify, and YouTube. The song quickly became viral, accumulating over 20 million views across all platforms [^50] https://www.nbcnews.com/pop-culture/viral-ai-powered-drake-weeknd-song-removed-streaming-services-rcna80098, before its removal by Universal Music Group, Drake’s recording label. In contrast, Canadian artist Grimes has extended an invitation to musicians to emulate her voice via AI for the creation of new musical pieces, stipulating that the lyrics should not be harmful. She advocates for the democratization of art and the abolition of copyright [^51] https://www.bbc.com/news/entertainment-arts-65385382. Additionally, Grimes has employed AI to design visual content for her LED backdrop at Coachella in 2024.
@@ 4.2 The future of AI technologies
A substantial amount of data is essential for training AI models to achieve high performance. Major companies such as Google, Meta, and NVIDIA, with their respective models—BERT, Segment Anything, and Canvas—dominate this space, leveraging their resources to gather data and train sophisticated models. However, Bloomberg released an article in November 2024 stating that OpenAI, Anthropic, and Google are experiencing slow growth in the performance of their AI models, with one of the key challenges being training data [^52] https://www.bloomberg.com/news/articles/2024-11-13/openai-google-and-anthropic-are-struggling-to-build-more-advanced-ai. Within the creative industries, a popular practice is to fine-tune these pre-trained models using techniques such as supervised learning with specific datasets, zero/few-shot learning, or domain adaptation.
@@
LLMs excel in applications involving complex tasks, advanced reasoning, data analysis, and understanding context. However, these models typically require high computational resources or cloud computing for operation and fine-tuning. A new trend emerging in parallel with LLMs is the development of Small language models (SLMs), such as Phi-3 by Microsoft [^53] https://azure.microsoft.com/en-us/blog/introducing-phi-3-redefining-whats-possible-with-slms/. SLMs offer promising solutions for regulated industries and sectors encountering scenarios where high-quality results are essential while keeping data on-premises. This potential is particularly evident in deploying more capable SLMs on smartphones and other mobile devices, allowing them to operate ‘at the edge’ without relying on cloud connectivity. The recently highly successful platforms, such as DeepSeek-V3 [376] and Qwen2.5-Max [377], are based on Mixture-of-Experts (MoE) models, which tackle complex problems by dividing them into simpler sub-tasks, each handled by a specialized “expert.”
@@
Despite the advancements in AI, the current models still struggle with tasks requiring planning ability and are prone to errors when encountering unexpected data. They may learn through reinforcement learning, but this process often identifies the best outcome as an anomaly rather than the norm. Yann LeCun, Professor at NYU and Chief AI Scientist at Meta, notes that while LLMs show a degree of comprehension in processing and generating text, their understanding lacks depth, leading to errors that defy common sense [^54] https://twitter.com/ylecun/status/1728496457601183865. He advocates for self-supervised learning as a pivotal future direction for AI, emphasizing its potential to derive insights from unlabeled data. Concurrently, Andrew Ng, Adjunct Professor at Stanford University and Founder of DeepLearning.AI, sees iterative AI agentic workflows [^55] https://www.youtube.com/watch?v=sal78ACtGTc as a key advancement for enhancing AI tool capabilities through an interactive approach by AI agents. These workflows utilize advanced algorithms and data processing techniques to automate complex tasks with simple data entry and validation to intricate decision-making processes.
@@
The increasing openness of code and datasets is seen as a catalyst for accelerating AI advancements, with major firms like Microsoft, Google, and Meta supporting open access technologies. However, the openness also introduces significant security risks, necessitating new regulatory measures to monitor models post-release, standardize documentation, and assess the safety of disclosing software code and training data.
@@
Finally, as mentioned in [2], the rapid advancement of these technologies has revolutionized cultural experiences, often referred to as ‘CreaTech’—the convergence of the creative and digital sectors [378]. These innovations not only reshape how people engage with art and creative works (e.g., through AR/VR) but also drive the evolution of the technologies themselves.
5 Concluding Remarks
This paper presents a comprehensive review of current AI technologies and their applications that have emerged in recent years. With generative AI, there is a rapid growth in AI usage in the creative sector, and these technologies have significantly advanced the state of the art across various creative applications, including content creation, information analysis, content enhancement, information extraction, and data compression.
Through these applications, Generative AI not only broadens creative possibilities but also reduces the manual labor and time traditionally associated with production pipeline, allowing for greater creative experimentation and quicker production cycles. As this technology advances, it promises to unlock even more sophisticated capabilities in the creative industry. However, artists and users need to adapt and learn to use these tools efficiently and effectively.

### v3

@@ 3.7.2 Performance and main challenges
Due to the lack of standard test conditions and limited model generalization within many existing image and video quality assessment models, deep compression methods are typically trained and benchmarked using different databases in conjunction with intra-database cross-validation. This can result in inconsistent evaluation results and conclusions. To enable a fair and meaningful comparison, various challenges and contests have been held for visual quality assessment. The Sixth Challenge on Learned Image Compression (CLIC) [336] associated with the Data Compression Conference 2024 is one of the latest examples which includes two quality assessment tracks for image and video compression. The best performer in the video quality assessment track achieves a Spearman Ranking Correlation Coefficient value of 0.825 [252], which is based on a ranking-inspired training methodology. Other notable challenges include the IEEE/CVF WACV 2023 HDR VQA Grand Challenge and the Video Super-Resolution Quality Assessment Challenge in ECCV 2024, which focus on high dynamic range and super-resolved content, respectively.
Although significant progress has been made in the past few years in visual quality assessment, including new models and training methodology, challenges remain, including limited model generalization and high computational complexity.
Another important use of quality metrics is as embedded loss functions for image and video processing optimization. This requires further capability and robustness, alongside complexity reduction, all topics to be addressed in future work.
4 Closing Thoughts and Future of AI in Creativity
This paper has presented a comprehensive review of current AI technologies and their applications that have emerged in recent years. Generative methods have driven a rapid growth in AI usage, particularly in the creative sector, significantly advancing the state of the art across various creative applications such as content creation, information extraction and analysis, content enhancement and data compression.
Through these applications, generative AI has not only broadened creative possibilities but has also reduced the manual effort and time traditionally associated with the production pipeline, allowing for greater creative experimentation and quicker production cycles. As this technology advances, it promises to unlock even more sophisticated capabilities in the creative industry. However, creative technologists, artists and other users must adapt, learn to use, and build these tools effectively and safely.
4.1 Challenges for AI in the Creative Sector
One of the primary challenges for artists engaging with modern generative AI and LLMs is the lack of consistent, controllable output. These models operate via stochastic sampling from high-dimensional latent spaces, meaning that identical prompts can yield different results across runs. This unpredictability makes it difficult for artists to achieve and iterate toward a precise creative vision. Although prompt engineering has emerged as a technique to guide model behavior, it requires technical knowledge and iterative refinement, which may not align with the intuitive or exploratory approaches common in artistic practice.
Moreover, there is a fundamental tension between the structured nature of current AI pipelines and the nonlinear, often improvisational workflows of creative disciplines. Many generative tools were originally designed for tasks like software development, content automation, or optimization [419], and are ill-suited for open-ended, exploratory creation. Artists typically work in cycles of ideation, experimentation, and revision—processes that demand fluid, real-time interaction and control, which existing AI systems struggle to support. These limitations point to a gap in current AI design: a need for systems that not only generate high-quality content but also adapt to the iterative, interpretive nature of artistic production. One possible approach addressed to these challenges is a shift toward top-down creative workflows, where artists define high-level concepts, themes, or goals via text prompts before refining specific outputs. This approach helps align AI-generated results with artistic intent, offering a degree of control over inherently stochastic systems.
Speaking at the World Government Summit in Dubai in 2024, [^56] https://blogs.nvidia.com/blog/world-governments-summit/ NVIDIA CEO Jensen Huang argued that with rapid advancements in AI, learning to code may become less essential for newcomers to the tech sector. He envisioned a future where traditional programming could be replaced by more intuitive AI-driven tools, thereby automating complex tasks and enhancing productivity—particularly for artists without coding expertise. While this perspective remains debated, it highlights the potential for AI to become more accessible within creative fields. However, AI-assisted coding tools are insufficient for creative practitioners, as artistic workflows tend to be unstructured and rely on domain-specific data. Artists must hence turn to techniques such as fine-tuning pre-trained models, few-shot learning, or domain adaptation—methods that are powerful yet typically inaccessible without machine learning expertise.
There is also broader concerns persist regarding the long-term impact of AI on the creative industries, particularly with the potential emergence of artificial general intelligence (AGI). Envisioned by organizations like OpenAI, DeepMind, and Anthropic, AGI could surpass human cognitive abilities, raising ethical and existential questions about the role of human agency in artistic expression.
@@ 4.2 Ethical Issues, Fakes and Bias
Advancements in generative AI, exemplified by models like Sora and Gemini 1.5 Pro, provoke ethical concerns and societal implications. These models, capable of generating highly realistic content, escalate the risk of misuse, through malicious deepfakes and misinformation. We are now in a situation where AI results transcend the uncanny valley, further complicating matters and challenging perceptions of authenticity. For example, the artist Miles Astray demonstrated that even authentic photographs could be mistaken for AI-generated images. His real photograph ‘F L A M I N G O N E’ won both the jury’s award and the people’s choice award in the AI category of the 1839 Awards. His aim wasto highlight the ethical dilemmas inherent in AI, suggesting that the benefits of discussing AI’s ethical implications could surpass the ethical concerns related to viewer deception [^57] https://www.milesastray.com/news/newsflash-reclaiming-the-brain.
While democratizing AI tools no-doubt presents opportunities to transform creative processes and workflows, it also necessitates robust regulatory frameworks to safeguard privacy and ownership. For example, deepfake technologies stimulate significant concerns about the spread of misinformation and other malicious uses. Efforts to detect and identify increasingly realistic deepfakes are thus as important as the generative methods used to produce them. These must however be accompanied by increased media literacy, and policies that address the ethical and legal implications.
@@
Diversity and representation is a key issue when using AI tools. Unified Concept Editing [54] has been proposed as a basis for image generation in digital mediums. This aims to ensure the production of safe content with diverse representation, reducing gender and racial biases. Hallucination in generative AI (the production of outputs that are not faithful representations of reality but instead contain imagined or unrealistic elements) are a further cause of concern. These undermine trust in AI processes and can be due to limitations in the training data, biases in the model architecture or imperfections in the optimization process. Hallucinations associated with LLMs are one of the issues highlighted by the UK Government [1], alongside bias, regurgitation of private data, difficulties with multi-step tasks and challenges in interpreting black-box processes.
@@
Governments across the world are increasingly expressing concerns about the challenges and uncertainties that generative AI technologies pose to rights holders and human creativity [2]. Generative AI presents substantial legal challenges, including the copyright status of AI-generated work and the intellectual property and copyright implications of the datasets used in training AI models. Viewpoints on this issue do however differ. For example, the track “Heart on My Sleeve,” penned by an (as yet unidentified) human author, featured AI-generated vocals that replicated the voices of Drake and The Weeknd. Released independently on April 4, 2023, it was accessible via streaming platforms including Apple Music, Spotify, and YouTube. The song quickly became viral, accumulating over 20 million views across all platforms [^58] https://www.nbcnews.com/pop-culture/viral-ai-powered-drake-weeknd-song-removed-streaming-services-rcna80098, prior to its removal by Universal Music Group, Drake’s recording label. In contrast, Canadian artist Grimes has extended an invitation to musicians to emulate her voice via AI for the creation of new musical pieces, stipulating that the lyrics should not be harmful. She has advocated for the democratization of art and the abolition of copyright [^59] https://www.bbc.com/news/entertainment-arts-65385382. Additionally, Grimes has employed AI to design visual content for her LED backdrop at Coachella in 2024.
Finally, the rapid development of AI technologies has also raised concerns about job displacement and the balance between automation and human participation in creative processes. Ensuring that AI augments, rather than undermines, human effort poses a significant challenge for developers and policymakers.
@@ 4.3 The future of AI technologies
Several key technological issues remain which need to be addressed if AI is to deliver its full potential. These in particular relate to training data, computational complexity and their depth of reasoning or planning, and are discussed below.
@@
A substantial amount of data is essential for training AI models in order to achieve high performance and good generalisation. Major companies such as Google, Meta, and NVIDIA, with their respective models: BERT, Segment Anything, and Canvas, dominate this space, benefiting from leveraged resources to gather data and process it to train sophisticated models. However, in November 2024, Bloomberg reported that OpenAI, Anthropic, and Google are all experiencing relatively slow growth in the performance of their AI models, with one of the key challenges being training data [^60] https://www.bloomberg.com/news/articles/2024-11-13/openai-google-and-anthropic-are-struggling-to-build-more-advanced-ai.
@@
LLMs excel in applications involving complex tasks, advanced reasoning, data analysis, and understanding context. However, these models typically require high computational resources or cloud computing for development, operation and fine-tuning. A new trend emerging in alongside LLMs is the development of Small language models (SLMs), such as Phi-3 by Microsoft [^61] https://azure.microsoft.com/en-us/blog/introducing-phi-3-redefining-whats-possible-with-slms/. SLMs offer promising solutions for regulated industries and sectors encountering scenarios where high-quality results are essential while keeping data ’on-site’. Their potential is particularly relevant when deploying more capable SLMs on smartphones and other mobile devices, allowing them to operate ‘at the edge’ without relying on cloud connectivity. Recent highly successful platforms, such as DeepSeek-V3 [420] and Qwen2.5-Max [421], are based on Mixture-of-Experts (MoE) models, which tackle complex problems by dividing them into simpler sub-tasks, each handled by a specialized “expert.”
@@
Despite the evident advancements in AI, current models still struggle with tasks requiring planning or deep reasoning and are prone to errors when encountering unexpected data. This, in turn, reduces the confidence of users and trust in the results. AI algorithms can learn through reinforcement learning, but this process often identifies the best outcome as an anomaly rather than the norm. Yann LeCun, Professor at NYU and Chief AI Scientist at Meta, noted that while LLMs show a degree of comprehension in processing and generating text, their understanding lacks depth, often leading to results that defy common sense [^62] https://twitter.com/ylecun/status/1728496457601183865. He advocates for self-supervised learning as a pivotal future direction for AI, emphasizing its potential to derive insights from unlabeled data. Concurrently, Andrew Ng, Adjunct Professor at Stanford University and Founder of DeepLearning.AI, sees iterative AI agentic workflows [^63] https://www.youtube.com/watch?v=sal78ACtGTc as a key advancement for enhancing AI tool capabilities through an interactive approach by AI agents. These workflows involve autonomous agents that interactively learn from experience, understand natural language, and execute tasks on behalf of users.
@@
The increasing openness of code and datasets is seen by many as a catalyst for accelerating AI advancements, with major firms like Microsoft, Google, and Meta supporting open access technologies. However, this openness also introduces security risks, necessitating new regulatory measures to monitor models post-release, to standardize documentation, and to assess the safety of software code and training data disclosure.
@@
Finally, as stated in [2], the rapid advancement of AI technologies has revolutionized cultural experiences, often referred to as ‘CreaTech’—the convergence of the creative and digital sectors [422]. Such innovations not only reshape how people engage with art and creative work (e.g., through AR/VR/MR) but also drive the evolution of the technologies themselves.